# references.bib

### v1

@@ Abstract
Multilevel Monte Carlo (MLMC) is a flexible and effective variance reduction technique for accelerating reliability assessments of complex power system. Recently, data-driven surrogate models have been proposed as lower-level models in the MLMC framework due to their high correlation and negligible execution time once trained. However, in resource adequacy assessments, pre-labeled datasets are typically unavailable. For large-scale systems, the efficiency gains from surrogate models are often offset by the substantial time required for labeling training data. Therefore, this paper introduces a speed metric that accounts for training time in evaluating MLMC efficiency. Considering the total time budget is limited, a vote-by-committee active learning approach is proposed to reduce the required labeling calls. A case study demonstrates that, within practical variance thresholds, active learning enables significantly improved MLMC efficiency with reduced training effort, compared to regular surrogate modelling approaches.
@@ I Introduction
In response to the increasing variability introduced by higher shares of renewable energy sources (RES) and electrified demand, probabilistic evaluations are widely implemented in power system operation and planning, assessing the impact of uncertainty on system reliability [spilger_uncertainty_2025]. In both administrative and research contexts, two widely used expectation-based metrics are Loss of Load Expectation (LOLE) [stephen_clarifying_2022] and Expected Energy Not Served (EENS) [evans_minimizing_2019].
Monte Carlo (MC) simulation is a commonly used technique, capable of evaluating the performance of complex systems under multiple sources of uncertainty. However, standard MC simulation suffers from slow convergence, particularly when estimating resource adequacy metrics such as LOLE [stephen_clarifying_2022], where typical standards are e.g., 3 event-hours in a year or 1 event-day in 10 years. Obtaining high-confidence reliability estimates therefore requires evaluating large number of scenarios. To address this, variance reduction techniques [liu_assessment_2021] have been developed to lower the estimation variance and thereby reduce the simulation count, but careful modifications to the simulation model are usually required.
Surrogate models, approximating system performance using fitted functions, have been proposed to bypass time-consuming simulations and thereby reduce the computational cost of scenario evaluation [li_surrogate_2019, gao_intelligent_2023]. However, the resulting uncertainty estimates are affected by the prediction errors inherent in surrogate models. Multilevel Monte Carlo (MLMC), as a flexible variance reduction approach, employs a hierarchy of correlated approximating models with different complexity [aslett_multilevel_2017]. Complex models are applied to correct the estimation bias made by reduced models, providing unbiased estimates with uncertainty confidence estimations. When well-trained, surrogate models produce highly correlated outputs with top-level simulations while requiring negligible execution time, making them particularly well suited as lower-level models in the MLMC framework [sharifnia_multilevel_2022, masoumi_deep-learning-enhanced_2024].
@@
To mitigate this issue while maintaining accuracy, the number of complex simulation calls should be minimized by maximizing their utility. Active learning offers a potential solution by selectively labeling only the most informative scenarios, so that each optimization call contributes maximally to improving the surrogate model’s accuracy [priesmann_artificial_2024].
@@
  • The MLMC speed metric from [tindemans_accelerating_2020] is adapted to include the impact of surrogate model training time.
@@ II-A MLMC framework for adequacy assessment
In contrast, MLMC employs a hierarchy of models $f_{1}(z),f_{2}(z),\cdots,f_{L-1}(z)$ with increasing fidelity and complexity to approximate $f_{L}(z)\equiv f(z)$. These reduced models are much easier to solve, but generally have an approximation error so that $\mathbb{E}\left[f_{l}(Z)\right]\neq\mathbb{E}\left[f_{L}(Z)\right]$ [tindemans_accelerating_2020]. Using $f_{0}\equiv 0$, the expectation-based metric $q$ can be decomposed into $L$ levels:
@@
The variance of the MLMC estimator with an optimal allocation of samples for a given simulation time budget $t_{sim}$ is given by [tindemans_accelerating_2020]
@@
Here, $\tau_{l}$ is the time for generating and evaluating the pair of samples at level $l$. The variance decreases with higher correlation between adjacent-level models. Furthermore, although using biased approximations with reduced models, MLMC provides an unbiased estimate of $q$, comparable to that of conventional MC methods, ensuring no loss of accuracy while saving computational resources [sharifnia_multilevel_2022].
@@ II-B Surrogate models for adequacy assessment
Surrogate models [li_surrogate_2019], also referred to as meta-models [priesmann_artificial_2024], have been widely applied to replace heavy simulations in energy system modeling [jiang_surrogate_2022, gao_intelligent_2023, masoumi_deep-learning-enhanced_2024]. To sufficiently capture the actual system behavior across all possible scenarios $Z$, the actual model $f(z)$ must be firstly called over a large number of sampled scenarios to generate a labeled training set. Surrogate models are trained with this training set and once trained, surrogate models generate outputs much more rapidly.
@@ II-C Speed measures for MLMC with surrogate models
where $\sigma_{\hat{Q}}$ denotes the variance of the estimator and $q$ is the targeted expectation value. The performance of a given MC algorithm, including MLMC, can be characterised by relating $c$ (lower is better) to total computation time $t$. This way, we can identify the computational ‘speed’ (linear in time) of an MC algorithm as [tindemans_accelerating_2020]
@@ III Active learning strategy for training surrogate models
The performance of data-driven surrogate models heavily depends on the labeled data used for training. This includes both the quantity of data samples and their informativeness, which contribute to the model’s generalization ability on unseen real-world scenarios [li_active_2024]. In the application of adequacy assessments, historical labeled dataset is often limited or entirely unavailable for training. Each data point must therefore be labeled through computationally expensive operation models. Suffering from the high computation burden, the number of samples that can be evaluated is thus constrained, making the effective selection of training samples crucial.
An active learning (AL) approach is therefore proposed to selectively label on the samples with high informativeness, maximizing the contribution of each data point in improving the generalization ability and therefore saving computation time for expensive labeling simulation [settles_active_2009].
@@
  • Vote-by-committee (disagreement-based) [bamdad_building_2020]: An ensemble of surrogate models is trained, e.g. through bootstrapping of the training data and randomizing initial conditions. Each model is used to predict outcomes for the unlabeled scenarios and the degree of disagreement (e.g. measured by the variance) is used to select points with output uncertainty.
  • Uncertainty sampling (confidence-based) [mahmood_active_2024]: For surrogate model types that explicitly provide confidence estimates, scenarios with low confidence are selected for evaluation.
@@ IV-A System description
To investigate the impact of surrogate model training time on MLMC performance when surrogate models are used as lower-level models, a case study of resource adequacy assessment is implemented on a single-node power system (also used in [sharifnia_multilevel_2022]). The system includes aggregated loads, wind turbines, 12 thermal generators, and 27 storage units operated under an EENS-minimizing dispatch policy [evans_minimizing_2019]. The exact model $f(z)$, based on simulation, evaluates the system’s loss-of-load (LOL) and energy not served (ENS) using a one-year generation margin time-series $z$ with 1 hour resolution:
@@ IV-B Surrogate model trained via active learning
Two random forest regressors, $\hat{f}^{LOL(day)}$ and $\hat{f}^{ENS(day)}$ are built to estimate daily LOL [h/day] and ENS [MWh/day] based on 24-hour [^1] We note that training a model on shorter traces will necessarily introduce a bias in the surrogate model, but this is corrected within the MLMC framework. margin time-series $z_{d}$. Random forests are widely applied ensemble models, naturally consisting of a committee of decision tree models for active learning, providing robust prediction while inexpensive to train [pham_predicting_2020]. The final estimates for the one-year LOL and ENS are computed by aggregating the prediction across 365 days:
@@
To evaluate the effectiveness of active learning in selecting informative samples for labeling, the vote-by-committee approach was implemented, as demonstrated on the left panel of Fig. 1. At the initial step, 730 daily margin time series were randomly generated and labeled using the exact model. Entering the active learning phase, given that storage units follows the EENS-minimizing dispatch policy, the ENS estimator was used to predict ENS values for a temporary pool of 3650 new daily margin traces. For each candidate trace, the standard deviation of predictions across individual decision trees in the ENS random forest was computed. The 91 days with the highest standard deviation (most disagreement) were then selected. Their true LOL and ENS values were obtained through simulation and added to the training set for the next iteration. For benchmarking, surrogate models were also trained using training sets constructed by random selecting daily margin traces.
All simulations were repeated for 10 times and the averaged values are presented in Section V. Code was written in Python and executed on a 14-inch MacBook Pro with an Apple M1 Pro processor (16 GB RAM). All code and data are available for download. [^2] https://github.com/RookieRQ/AL-enhanced-Surrogate_MLMC-Adequacy_Assessment
@@ V Results
Exact model N/A 0 2004 $1.753\pm 0.146$ 0.118 N/A $2417.2\pm 209.0$ 0.067 N/A
AL (5 rounds) $1185$ 13.53 2004 $1.749\pm 0.058$ 0.829 1.86 $2454.6\pm 136.3$ 0.296 1.18
AL (10 rounds) $1640$ 30.65 2004 $1.755\pm 0.034$ 1.462 32.77 $2442.7\pm 87.5$ 0.526 11.59
AL (20 rounds) $2550$ 72.67 1999 $1.739\pm 0.028$ 2.009 225.63 $2369.0\pm 60.6$ 0.891 53.95
Random surrogate 3285 14.58 1996 $1.724\pm 0.050$ 0.694 2.07 $2393.2\pm 140.8$ 0.192 1.51
7300 33.59 2004 $1.743\pm 0.039$ 1.158 62.02 $2396.6\pm 87.4$ 0.473 6.14
14600 72.24 2000 $1.742\pm 0.037$ 1.232 745.14 $2411.6\pm 89.6$ 0.468 Invalid
@@ VI Conclusions and future work
[…]

### v2

# MLMC-based Resource Adequacy Assessment with Active Learning Trained Surrogate Models

@@ Abstract
Multilevel Monte Carlo (MLMC) is a flexible and effective variance reduction technique for accelerating reliability assessments of complex power system. Recently, data-driven surrogate models have been proposed as lower-level models in the MLMC framework due to their high correlation and negligible execution time once trained. However, in resource adequacy assessments, pre-labeled datasets are typically unavailable. For large-scale systems, the efficiency gains from surrogate models are often offset by the substantial time required for labeling training data. Therefore, this paper introduces a speed metric that accounts for training time in evaluating MLMC efficiency. Considering the total time budget is limited, a vote-by-committee active learning approach is proposed to reduce the required labeling calls. A case study demonstrates that, within a given computational budget, active learning in combination with MLMC can result in a substantial reduction variance.
@@ I Introduction
In response to the increasing variability introduced by higher shares of renewable energy sources (RES) and electrified demand, probabilistic evaluations are widely implemented in power system operation and planning, assessing the impact of uncertainty on system reliability [1]. In both administrative and research contexts, two widely used expectation-based metrics are Loss of Load Expectation (LOLE) [2] and Expected Energy Not Served (EENS) [3].
Monte Carlo (MC) simulation is a commonly used technique, capable of evaluating the performance of complex systems under multiple sources of uncertainty. However, standard MC simulation suffers from slow convergence, particularly when estimating resource adequacy metrics such as LOLE [2], where typical standards are e.g., 3 event-hours in a year or 1 event-day in 10 years. Obtaining high-confidence reliability estimates therefore requires evaluating large number of scenarios. To address this, variance reduction techniques such as importance sampling [4] have been developed to alter the sampling strategy to focus on high-impact scenarios samples, thereby reducing the scenario count, but there is a tight coupling between the simulation model and the available sampling strategies.
Surrogate models, approximating system performance using fitted functions, have been proposed to bypass time-consuming simulations and thereby reduce the computational cost of scenario evaluation [5, 6]. However, the resulting uncertainty estimates are affected by the prediction errors inherent in surrogate models. Multilevel Monte Carlo (MLMC), as a flexible variance reduction approach, employs a hierarchy of correlated approximating models with different complexity [7]. Complex models are applied to correct the estimation bias made by reduced models, providing unbiased estimates with uncertainty confidence estimations. When well-trained, surrogate models produce highly correlated outputs with top-level simulations while requiring negligible execution time, making them particularly well suited as lower-level models in the MLMC framework [8, 9].
@@
To mitigate this issue while maintaining accuracy, the number of complex simulation calls should be minimized by maximizing their utility. Active learning offers a potential solution by selectively labeling only the most informative scenarios, so that each optimization call contributes maximally to improving the surrogate model’s accuracy [10].
@@
  • The MLMC speed metric from [11] is adapted to include the impact of surrogate model training time.
@@ II-A MLMC framework for adequacy assessment
In contrast, MLMC employs a hierarchy of models $f_{1}(z),f_{2}(z),\cdots,f_{L-1}(z)$ with increasing fidelity and complexity to approximate $f_{L}(z)\equiv f(z)$. These reduced models are much easier to solve, but generally have an approximation error so that $\mathbb{E}\left[f_{l}(Z)\right]\neq\mathbb{E}\left[f_{L}(Z)\right]$ [11]. Using $f_{0}\equiv 0$, the expectation-based metric $q$ can be decomposed into $L$ levels:
@@
The variance of the MLMC estimator with an optimal allocation of samples for a given simulation time budget $t_{sim}$ is given by [11]
@@
Here, $\tau_{l}$ is the time for generating and evaluating the pair of samples at level $l$. The variance decreases with higher correlation between adjacent-level models. Furthermore, although using biased approximations with reduced models, MLMC provides an unbiased estimate of $q$, comparable to that of conventional MC methods, ensuring no loss of accuracy while saving computational resources [8].
@@ II-B Surrogate models for adequacy assessment
Surrogate models [5], also referred to as meta-models [10], have been widely applied to replace heavy simulations in energy system modeling [12, 6, 9]. To sufficiently capture the actual system behavior across all possible scenarios $Z$, the actual model $f(z)$ must be firstly called over a large number of sampled scenarios to generate a labeled training set. Surrogate models are trained with this training set and once trained, surrogate models generate outputs much more rapidly.
@@ II-C Speed measures for MLMC with surrogate models
where $\sigma_{\hat{Q}}$ denotes the variance of the estimator and $q$ is the targeted expectation value. The performance of a given MC algorithm, including MLMC, can be characterised by relating $c$ (lower is better) to total computation time $t$. This way, we can identify the computational ‘speed’ (linear in time) of an MC algorithm as [11]
@@ III Active learning strategy for training surrogate models
The performance of data-driven surrogate models heavily depends on the labeled data used for training. This includes both the quantity of data samples and their informativeness, which contribute to the model’s generalization ability on unseen real-world scenarios [13]. In the application of adequacy assessments, historical labeled dataset is often limited or entirely unavailable for training. Each data point must therefore be labeled through computationally expensive operation models. Suffering from the high computation burden, the number of samples that can be evaluated is thus constrained, making the effective selection of training samples crucial.
An active learning (AL) approach is therefore proposed to selectively label on the samples with high informativeness, maximizing the contribution of each data point in improving the generalization ability and therefore saving computation time for expensive labeling simulation [14].
@@
  • Vote-by-committee (disagreement-based) [15]: An ensemble of surrogate models is trained, e.g. through bootstrapping of the training data and randomizing initial conditions. Each model is used to predict outcomes for the unlabeled scenarios and the degree of disagreement (e.g. measured by the variance) is used to select points with output uncertainty.
  • Uncertainty sampling (confidence-based) [16]: For surrogate model types that explicitly provide confidence estimates, scenarios with low confidence are selected for evaluation.
@@ IV-A System description
To investigate the impact of surrogate model training time on MLMC performance when surrogate models are used as lower-level models, a case study of resource adequacy assessment is implemented on a single-node power system (also used in [8]). The system includes aggregated loads, wind turbines, 12 thermal generators, and 27 storage units operated under an EENS-minimizing dispatch policy [3]. The exact model $f(z)$, based on simulation, evaluates the system’s loss-of-load (LOL) and energy not served (ENS) using a one-year generation margin time-series $z$ with 1 hour resolution:
@@ IV-B Surrogate model trained via active learning
Two random forest regressors, $\hat{f}^{LOL(day)}$ and $\hat{f}^{ENS(day)}$ are built to estimate daily LOL [h/day] and ENS [MWh/day] based on 24-hour [^1] We note that training a model on shorter traces will necessarily introduce a bias in the surrogate model, but this is corrected within the MLMC framework. margin time-series $z_{d}$. Random forests are widely applied ensemble models, naturally consisting of a committee of decision tree models for active learning, providing robust prediction while inexpensive to train [17]. The final estimates for the one-year LOL and ENS are computed by aggregating the prediction across 365 days:
@@
The vote-by-committee approach was implemented as the active learning strategy for selecting informative samples, as demonstrated on the left panel of Fig. 1. At the initial step, 730 daily margin time series were randomly generated and labeled using the exact model. Entering the active learning phase, given that storage units follows the EENS-minimizing dispatch policy, the ENS estimator was used to predict ENS values for a temporary pool of 3650 new daily margin traces. For each candidate trace, the standard deviation of predictions across individual decision trees in the ENS random forest was computed. The 91 days with the highest standard deviation (most disagreement) were then selected. Their true LOL and ENS values were obtained through simulation and added to the training set for the next iteration. For benchmarking, surrogate models were also trained using training sets constructed by random selecting daily margin traces.
This study uses a basic setting of the random forest regressor, with 100 decision trees and vote-by-committee active learning. While these settings are sufficient to demonstrate the concept, surrogate model performance could be enhanced through hyperparameter tuning and exploration of alternative active learning methods. Moreover, the efficiency of the surrogate-assisted MLMC framework could be further improved through the use of importance sampling strategies.
All simulations were repeated for 10 times and the averaged values are presented in Section V. Code was written in Python and executed on a 14-inch MacBook Pro with an Apple M1 Pro processor (16 GB RAM). All code and data are available for download. [^2] Zenodo, May 27, 2025. doi: 10.5281/zenodo.16600719.
@@ V Results
Exact model N/A 0 2004 1.753±0.1461.753$\pm$ 0.1461.753 ± 0.146 0.118 N/A 2417.2±209.02417.2$\pm$ 209.02417.2 ± 209.0 0.067 N/A
AL (5 rounds) 118511851185 13.53 2004 1.749±0.0581.749$\pm$ 0.0581.749 ± 0.058 0.829 1.86 2454.6±136.32454.6$\pm$ 136.32454.6 ± 136.3 0.296 1.18
AL (10 rounds) 164016401640 30.65 2004 1.755±0.0341.755$\pm$ 0.0341.755 ± 0.034 1.462 32.77 2442.7±87.52442.7$\pm$ 87.52442.7 ± 87.5 0.526 11.59
AL (20 rounds) 255025502550 72.67 1999 1.739±0.0281.739$\pm$ 0.0281.739 ± 0.028 2.009 225.63 2369.0±60.62369.0$\pm$ 60.62369.0 ± 60.6 0.891 53.95
Random surrogate 3285 14.58 1996 1.724±0.0501.724$\pm$ 0.0501.724 ± 0.050 0.694 2.07 2393.2±140.82393.2$\pm$ 140.82393.2 ± 140.8 0.192 1.51
7300 33.59 2004 1.743±0.0391.743$\pm$ 0.0391.743 ± 0.039 1.158 62.02 2396.6±87.42396.6$\pm$ 87.42396.6 ± 87.4 0.473 6.14
14600 72.24 2000 1.742±0.0371.742$\pm$ 0.0371.742 ± 0.037 1.232 745.14 2411.6±89.62411.6$\pm$ 89.62411.6 ± 89.6 0.468 Invalid
@@ VI Conclusions and future work
This paper investigates the impact of surrogate model training time on the overall MLMC performance, explicitly considering the total time budget. A modified speed metric is proposed, showing that the asymptotic speedup achieved through surrogate models is offset by a factor dependent on both the training time and the total time budget. To better utilize the training effort, a vote-by-committee active learning strategy using random forest regression is employed to reduce the number of simulations for labeling. One case study of resource adequacy assessment shows that active learning leads to improved surrogate accuracy and greater MLMC efficiency compared to random sampling. Future work will focus on extending studies to large-scale applications, and further optimizing the labeling process by leveraging information from the high-fidelity model during MLMC execution.